OpenStudy (anonymous):

solve 8+ sqrt 2x≤5 the sqrt is only over 2x

OpenStudy (anonymous):

@jim_thompson5910 I got x≤9/2 for my anwer is that right?

jimthompson5910 (jim_thompson5910):

\[\Large 8+\sqrt{2x} \le 5\] \[\Large \sqrt{2x} \le 5-8\] \[\Large \sqrt{2x} \le -3\] is that inequality ever true for any x value?

OpenStudy (anonymous):

so Im correct right?

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

is \[\Large \sqrt{2x} \le -3\] ever true?

OpenStudy (anonymous):

no

jimthompson5910 (jim_thompson5910):

so there are no solutions

OpenStudy (anonymous):

ohh ok